### v1

# Catalog of Mock Stellar Streams in Milky Way-like Galaxies

Abstract
Dynamically cold stellar streams from tidally dissolved globular clusters (GCs) serve as excellent tools to measure the Galactic mass distribution and show promise to probe the nature of dark matter. For successful application of these tools to observations, it is essential to have an accurate model of stellar stream properties on the Galactic scale. To this end we produce a mock catalog of stellar streams in four simulated Milky Way-like galaxies. We build the catalog with three main components: a model for the formation and disruption of globular clusters based on cosmological simulations, time-dependent potentials constructed with basis function expansions for integrating stream orbits, and an improved particle spray algorithm for efficient generation of stellar streams. We find that the observable widths and lengths of mock streams as a function of galactocentric radius are well described by power-laws for streams beyond 10 kpc. We generate mock photometry for Gaia, LSST, and Roman, and find that the latter two surveys will increase the number of observable stars in GC stellar streams by several orders of magnitude. Our full catalog, containing stream populations across four different galaxy realizations, is publicly available and can be used to study stream population statistics and to calibrate models which use stellar streams to understand our Galaxy.
keywords:
Stellar streams — Globular star clusters — Stellar dynamics — Galaxy dynamics
1 Introduction
Globular clusters (GCs) are some of the best available tools for understanding the assembly history and evolution of the Milky Way galaxy (MW). Since most clusters are observed to have an age $\sim 10$ Gyr (Krauss & Chaboyer, 2003; Dotter et al., 2011), they have been orbiting our Galaxy for at least a few Gyr if accreted onto the MW (ex-situ formed clusters) or longer if born in the central part of the MW (in-situ formed clusters). As GCs orbit the Galaxy, they are subject to tidal forces that cause some stars to leave the cluster and form an unbound stellar stream. Recently, the Gaia mission (Gaia Collaboration et al., 2018, 2023) has significantly expanded our knowledge of MW streams (Bonaca & Price-Whelan, 2025), of which $\sim$80 are thought to originate from GCs (Mateu, 2023). Since the velocity dispersion of unbound GC stars is smaller than the typical orbital velocity around the MW, GC streams are dynamically cold, retaining signatures of their tidal histories for up to several Gyr.
One of the most exciting applications which shows promise to be better understood through stellar streams is constraining the mass distribution within the Galaxy (Johnston et al., 1999; Helmi et al., 1999; Koposov et al., 2010; Sanders & Binney, 2013; Bonaca & Hogg, 2018; Reino et al., 2021; Ibata et al., 2024; Nibauer & Bonaca, 2025). This is typically done by assuming an analytic form of the gravitational potential and then varying the potential’s free parameters and/or symmetry constraints to identify which model can best recreate a given stream or select group of streams. In addition, Nibauer & Bonaca (2025) showed that the GD-1 stream (Grillmair & Dionatos, 2006) can be used to constrain the potential by direct differentiation of the stream track in phase space, without first having to assume an analytic form for the potential. Stream modeling can also be used to infer properties of satellite galaxies interacting with the MW. For example, Shipp et al. (2021) used measurements from the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016), the Southern Stellar Stream Spectroscopic Survey ($S^{5}$; Li et al., 2019), and Gaia to estimate the mass of the Large Magellanic Cloud (LMC) by modeling its perturbations on different stellar streams, obtaining a value of $1.88^{+0.35}_{-0.40}\times 10^{11}\,{\mathrm{M}_{\odot}}$. Vasiliev et al. (2021) performed a similar type of analysis using restricted N-body modeling to recreate Gaia measurements of the Sagittarius stream (Ibata et al., 2001) in the presence of the LMC, computing a MW virial mass of $(9.0\pm{1.3})\times 10^{11}\,{\mathrm{M}_{\odot}}$.
The density of stars within streams may exhibit variations caused by gravitational perturbations from substructure in the Galactic potential, such as a tilted stellar disk (Nibauer et al., 2024), giant molecular clouds (Amorisco et al., 2016), the Galactic bar (Pearson et al., 2017; Bonaca et al., 2020), and dark matter subhaloes (Johnston et al., 2002; Ibata et al., 2002; Carlberg, 2012; Erkal et al., 2016; Banik et al., 2018). Disentangling the baryonic perturbations from those potentially caused by dark matter (DM) is key in the hope of using stellar streams to constrain the lower end of the DM subhalo mass function. Detecting streams at large galactocentric radii will be extremely important for this purpose. A key effort in this direction is to model streams beyond the stellar disk ($r_{\rm gal}\sim 15\rm\penalty 10000\ kpc$), both in terms of physical properties and prospects for observing them.
Many published works model streams via the particle-spray method, which releases stream particles near the Lagrange points at frequent timesteps with small offsets in their positions and velocities (Varghese et al., 2011; Lane et al., 2012; Küpper et al., 2012; Bonaca et al., 2014; Gibbons et al., 2014; Fardal et al., 2015; Roberts et al., 2025). Integrating the orbits of these particles to the present day produces an elongated stream profile, generally consistent with the observations, providing a reasonably accurate method which is significantly faster than full N-body modeling methods. Recently, Chen et al. (2025b) developed an improved particle-spray algorithm where the phase-space offsets of released particles follow a multivariate Gaussian distribution. The new method matches the morphology and action space distribution of streams from N-body simulations with an error of only 10%. This method allows computationally efficient modeling of galactic stream populations with greater detail than previously possible.
Previous work has shown that due to the slow rate at which streams phase-mix with the Galactic stellar population, there is likely to be a large population of GC streams that remain distinguishable from field stars (Sandford et al., 2017; Carlberg, 2018), with many of them still undetected by the current observational facilities. In particular, Pearson et al. (2024) modeled the evolution of streams in a static MW potential using the Fardal et al. (2015) particle-spray method and found that there may be up to $\sim 1000$ coherent GC stellar streams in the MW. Of these, more than $100$ coherent GC stellar streams are predicted to be beyond $15\penalty 10000\ \rm kpc$ from the Galactic center, which would be observable by the LSST survey on the Vera C. Rubin Observatory with 10-year stacking data ($g<27.5$; LSST Science Collaboration et al., 2009).
In this paper, we build upon the work of Pearson et al. (2024) and generate mock GC stream populations in four simulated MW-like galaxies with realistic time-evolving potentials. We use the Chen et al. (2025b) particle spray algorithm to integrate the orbits of unbound stars for several Gyr. We analyze overall properties of the final stream populations and study how facilities such as LSST and the future Nancy Grace Roman Space Telescope (Roman; Spergel et al., 2013, 2015) will expand our ability to detect Galactic GC streams. We also make our stream catalog publicly available to be used as a tool for analyzing expected trends in the stream populations, and for calibration of models which attempt to understand the Galaxy through the use of GC streams.
The paper is organized as follows: in Section 2 we describe components of the model used to generate the catalog and introduce the data structure of the catalog itself. In Section 3 we describe properties of the streams in the catalog and evaluate photometric predictions for observing the streams. We discuss these results in further detail and provide more possible use cases of the catalog in Section 4. We conclude with a summary of our results in Section 5.
2 Model Setup
In this section, we describe the details of the modeling used to generate our mock stream catalog. The key components are the initialization of the GC population, dynamical evolution of the GCs, and the generation of streams via the particle spray method.
2.1 Globular Cluster Evolution
The modeling of GC evolution requires a detailed understanding of not only the birth of clusters, both in-situ and ex-situ, but also their mass loss over time through stellar evolution and tidal disruption. For these purposes, we employ the GC evolution model by Chen & Gnedin (2022), which is further developed in Chen & Gnedin (2024a). The model post-processes snapshots of a selected cosmological simulation and creates catalogs of the GC population, keeping track of cluster formation and evolution. The model produces key observables including mass, metallically, positions, and velocities of individual GCs. The model has been applied to a wide range of galaxies in cosmological simulations and can successfully match the observed properties of the Galactic GC system.
A full description of the most recent version of the model is given in Chen & Gnedin (2024a). The three main features are: cluster formation, particle assignment for orbit tracking, and mass loss through tidal disruption along the GC’s orbit. GC formation is triggered when the host galaxy mass accretion rate exceeds a specified threshold, which is an adjustable parameter of the model. The total mass of the cluster population formed at a given snapshot epoch is proportional to the cold gas mass of the parent galaxy, following Kravtsov & Gnedin (2005). Given this population, the individual masses of GCs are sampled from a Schechter (1976) initial cluster mass function with a lower bound of $10^{4}\,{\mathrm{M}_{\odot}}$ and an exponential cutoff of $10^{7}\,{\mathrm{M}_{\odot}}$. To obtain kinematic information for the clusters at this epoch, the model assigns clusters to collisionless particles in the background cosmological simulation and uses their positions and velocities. The mass evolution takes into account both stellar evolution and tidal disruption along the cluster orbit. Once a cluster is initialized with mass $M_{0}$, the model treats the mass-loss due to stellar evolution as instantaneous:
$M_{\rm i}=\mu_{\rm sev}M_{0}$ (1)
where the remaining fraction $\mu_{\rm sev}$ is taken to be $0.55$ (Gieles & Gnedin, 2023). This assumption is justified because the timescale for stellar evolution of high mass stars (tens of Myr) is much shorter than the typical age of GCs. Here $M_{\rm i}$ represents the initial mass of the cluster after stellar evolution but before any tidal disruption.
The stream is populated by stars unbound from the cluster by tidal forces. The most recent prescription for tidal mass-loss outlined in Chen & Gnedin (2024a) is given by:
$\frac{dM}{dt}=-2\times 10^{4}\frac{M_{\odot}}{\text{Gyr}}\left(\frac{M_{\rm i}}{2\times 10^{5}M_{\odot}}\right)^{\frac{1}{3}}\left(\frac{M(t)}{M_{\rm i}}\right)^{-\frac{1}{3}}\left(\frac{\Omega_{\mathrm{tid}}(t)}{150\,\text{Gyr}^{-1}}\right)$ (2)
where $\Omega_{\mathrm{tid}}(t)$ is the effective tidal frequency, approximated via the first and third eigenvalues ($\lambda_{1},\lambda_{3}$) of the tidal tensor:
$\Omega_{\mathrm{tid}}(t)^{2}=\lambda_{1}-\lambda_{3}$ (3)
where the tidal tensor is composed of the second derivatives of the Galactic potential $\Phi$. In Cartesian coordinates, it takes the form:
$\mathrm{T}_{ij}=-\frac{\partial^{2}\Phi}{\partial x^{i}\partial x^{j}}$ (4)
The model calculates the tidal tensor using a second order finite difference method for snapshots where the potential for the host galaxy is provided. However, this only permits the masses of the GCs to be updated at every full snapshot. To improve the time resolution, we use a combination of the GC masses computed by the Chen & Gnedin (2024a) model as well as a higher-resolution calculation of the mass-loss rate outlined in Section 2.3.
We note that by default, the GC evolution model calculates the mass of each GC at every snapshot, but only outputs the mass to the user at the present day. To address this for the purposes of this work, we make a small modification of the code to have it output the cluster masses at every snapshot. Due to this modification, we re-run the model on the four cosmological halos utilized in this work. Both re-running the model as well as utilizing a different mass-loss calculation method results in our GCs being slightly different from the GCs in the published catalogs of Chen & Gnedin (2024a). These differences are explained in detail in Appendix C, but we stress that overall our GCs are still tailored to match the MW system in terms of the mass, metallicity, and radial distributions.
2.2 Background Cosmological Simulations
To generate a realistic model for an initial GC population in MW-like galaxies, we utilize two halos from the Illustris TNG50-1 (TNG) cosmological simulation (Nelson et al., 2019), and two halos from the Latte suite of FIRE-2 (FIRE) simulations (Wetzel et al., 2016). The TNG galaxies have subhalo IDs 523889 and 519311 in the TNG halo catalog, and the FIRE halos are m12i and m12w. Chen & Gnedin (2024b) identified these four halos to be a close match to the Milky Way for the purpose of applying their GC post-processing model. The criteria used for determining suitable galaxies in Chen & Gnedin (2024b) include having a present mass and circular velocity similar to that observed for the MW, no major merger (defined as having a mass ratio greater than 1:4) in the last 10 Gyr, one major merger between 10-12 Gyr ago representing the Gaia-Sausage/Enceladus merger (Belokurov et al., 2018; Helmi et al., 2018), and to have formed 25-35% of its stellar mass by 10 Gyr ago (Leitner, 2012). Our selected halos satisfy all these requirements and are suitable candidates to approximate the present day stream distribution of the MW.
TNG and FIRE provide individual snapshots of halo particles, separated by $10^{8}-10^{9}$ yr, which by itself is not suitable for computing the orbits of millions of stream stars for several Gyr. For the actual computation of stream orbits, we construct analytical basis function expansion (BFE) potentials based on the TNG and FIRE snapshots using the agama software (Vasiliev, 2019). We construct the analytic time-dependent potentials in the following way: for each snapshot, we fit a BFE potential to the snapshot given the positions and masses of the stellar, gas, and dark matter particles. The dark matter particles are fit with a standard multipole expansion of spherical harmonics with $l_{\text{max}}$ taken to be $12$. The stellar and gas particles are fit with an azimuthal Fourier harmonic expansion with $m_{\text{max}}=6$. The method for determining these parameters and further details are outlined in Appendix A. We construct the time-variable potentials by linearly interpolating the expansion coefficients between the snapshots. For each simulation the overall time span is taken to be $z=0.3$ to the present day (about 3.5 Gyr). For the TNG halos, this corresponds to every snapshot starting from 78 to 99 (present day), which are spaced by intervals ranging from 120235120-235 Myr for TNG’s cosmological parameters. For FIRE, we use the same snapshots with full particle data as in Chen & Gnedin (2024b) from $z=0.3$ to present, corresponding to snapshots 446, 486, 534, and 590-600. The spacing between these snapshots respectively is 987 Myr, 1140 Myr, 1280 Myr, and 2.22 Myr for the remaining ten.
For the FIRE halo m12i, we find that the principal axis of the disk tilts by $24^{\circ}$ over this time span, and therefore we modify the above prescription for this halo only and model its potential just with multipole expansion. This is because all orbits are computed in the galactocentric frame of the halo at the last snapshot, but the agama Fourier harmonic expansion fitting procedure requires the disk to be aligned with the XYX-Y plane to work properly. Since multipole expansion is not as effective at modeling flattened systems, this results in our disk potential for m12i not being as accurate as in the other three halos. However, we find that fitting one of our other galaxies with only multipoles results in about 2% average difference in the orbits compared to using both BFE functions. Since this difference is small, we still proceed to use the m12i halo, but users of the catalog should be aware of this if they choose to use the m12i streams.
Instead of integrating orbits of particles immediately from cluster birth (or from time of accretion onto the MW), we select a lookback time to begin tracking the orbits of tidally stripped stars from each cluster. This is because phase mixing causes the orbits of evaporated stars to deviate significantly from the progenitor with time, limiting the duration of coherent streams on most orbits to a few Gyr. Moreover, Chen et al. (2025a) showed that the duration of most observed GC stellar streams are $\leq 1\rm\penalty 10000\ Gyr$ (although some streams are likely to be much older – see Erkal et al. 2016). It is thus unnecessary to include these early-released particles in our catalog. Based on these arguments we take our initial snapshot for orbit tracking of evaporated particles to be $z=0.3$.
We stress that the main benefit of our time-dependent potential is that it captures the mass evolution of the galaxies over the course of the GC orbits. For each halo there is significant mass growth over $3.5$ Gyr and therefore this effect is important. In principle, a time dependent potential could capture other aspects of a dynamic Galactic potential on stellar streams, such as the spiral arms and the galactic bar. However, the cadence of the simulation snapshots is insufficient to accurately capture these effects in the smoothed BFE potential.
2.3 Stellar Stream Generation
In this section, we describe the complete process for stream generation, given the Galactic potential and our initial GC conditions from the model of Chen & Gnedin (2024a). The distinct inputs to the model are as follows:
  • The Galactic potential $\Phi(\textbf{x},t)$
  • The GC masses at birth and at $z=0.3$
  • Positions and velocities of GCs at $z=0.3$
  • The cluster ages (at present day) and metallicities
We use the Galactic potentials as described above in Sec. 2.2. Every other item is supplied from the GC formation model, which we treat as input for generating our streams. Our stream generation has three main components: calculation of the cluster mass-loss rate, selection of released stream stars, and orbit integration of the stream stars.
2.3.1 Mass Loss Calculation
The overall mass-loss calculation is in two parts. From cluster birth until $z=0.3$, we use the mass loss given by the GC formation model. Then beginning at $z=0.3$, we numerically solve Equation 2 given the tidal tensor along cluster orbit computed with agama, to calculate the amount of mass lost by the cluster from $z=0.3$ until the present day. While agama uses an adaptive timestep for orbit integration, we interpolate the solution such that the mass history and tidal tensor components are provided at 1000 uniformly spaced intervals over the integration time, giving us both quantities with a time resolution of $3.5$ Myr. Figure 12 in Appendix A shows the effective tidal frequency as a function of time at the provided resolution for an example cluster in one of the TNG halos. There are only four full TNG snapshots with potential information in our chosen time of integration, which miss significant variation of the tidal force. The agama framework is much more accurate for detailed calculations of the mass loss.
2.3.2 Stream Realization
Once calculated, the mass-loss histories are passed onto the next step, where we initialize a stellar population for each cluster and compute the number and masses of stream stars to be released at each timestep. First, individual cluster stars are sampled according to a Kroupa (2001) IMF with the initial cluster mass given by the GC model. For the lower and upper bounds of the IMF, we use MIST isochrones (Choi et al., 2016; Dotter, 2016) to calculate the theoretical lowest and highest mass stars that would still be luminous today given the cluster’s age and metallicity. These values vary from cluster to cluster but are generally around $0.1\,{\mathrm{M}_{\odot}}$ for the lower bound and between $0.8-0.9\,{\mathrm{M}_{\odot}}$ for the upper bound. Stars with higher masses than this are already removed from the cluster mass accounting by stellar evolution in §2.1. We then evolve the mass function by removing stars until the sum of the remaining stellar mass is equal to the cluster mass at $z=0.3$. Since models of star cluster evolution (Baumgardt & Makino, 2003) show that low-mass stars tend to evaporate preferentially compared to high-mass stars, we remove stars probabilistically such that low-mass stars are more likely to be evaporated. Specifically, we treat the probability that any given star be removed as proportional to the inverse square root of its mass $m$:
$p_{\rm remove}\propto m^{-1/2}$ (5)
based on energy equipartition arguments. We continuously remove stars and re-normalize the probabilities until the sum of the remaining stellar masses is equal to the mass of the cluster at $z=0.3$. The remaining pool of stars are considered to be the bound cluster stars before generating the stream.
To further assess what effect the evaporation probability has on the stream profile, we selected a cluster and generated its stream twice, once with the adopted ejection probability $p\propto m^{-1/2}$ and once where every star is equally likely to be ejected independent of mass. After running the model, we binned the stars in the stream according to mass such that there were approximately 500 stars in each bin, and looked for variations between the two models as a function of mass. We found none significant. For example, the width of the stream, which we quantify with the standard deviation along the short axis of the stream, $\sigma_{\phi_{2}}$, shows little dependence on stellar mass: the best-fit slope of the linear regression was $0.16\pm 0.015$ for the stream with weighted ejection probability, and $0.010\pm 0.018$ for the stream with uniform ejection probability. Both of these slope are small, and we conclude that the evaporation probability has no significant effect on the structure of the stream and proceed to use it in our model to represent observations that older parts of streams have lower mass stars (Balbinot & Gieles, 2018).
Figure 1: Evolution of the mass function over time for a $2\times 10^{4}\,{\mathrm{M}_{\odot}}$ cluster on a Pal 5-like orbit with stellar ejection probability $p\propto m^{-1/2}$ and assumed metallicity of $\mathrm{[Fe/H]}=-1.56$ (the observed metallicity of Pal 5). The cluster becomes less dominated by low-mass stars gradually over time because they are more likely to be ejected than high-mass stars.
Figure 2: All sky map of the final collection of disrupted features for the TNG halo 523889 with $g<24.5$. Color indicates the metallicity of the original star cluster. Positions and apparent magnitudes of stars are calculated with the standard astropy convention where the Sun is placed at $(-8\rm\penalty 10000\ kpc,0\penalty 10000\ kpc,20.8\penalty 10000\ pc)$ in our galactocentric reference frame.
Analytically, the number of stars unbound from the cluster $N_{\rm stars}$ will depend on the mass-loss rate along the cluster’s orbit and the remaining mass function of stars in the cluster. This can be expressed as:
$N_{\rm stars}(t)=\int^{t}\frac{\dot{M}(t^{\prime})}{m_{\rm av}(t^{\prime})}\penalty 10000\ dt^{\prime}$ (6)
where $\dot{M}(t)$ is the mass loss rate along the progenitor orbit and $m_{\rm av}(t)$ quantifies the average mass of ejected stars at time t. In terms of the ejection probability for a given star $p(m,t)$ (time-dependent because the normalization factor decreases as the number of stars in the cluster decreases) and the mass function of the cluster $\phi(m,t)$, $m_{\rm av}$ is equal to
$m_{\rm av}(t)=\frac{\int p(m,t)\,m\,\phi(m,t)\,dm}{\int p(m,t)\,\phi(m,t)\,dm}$ (7)
Figure 1 shows how the mass function of an example cluster in our model changes over time given the ejection probability of the individual stars. We take a cluster with initial mass $2\times 10^{4}\,{\mathrm{M}_{\odot}}$, a Kroupa IMF with lower and upper bounds of $0.1-1.1\,{\mathrm{M}_{\odot}}$, similar apocenter and pericenter parameters as Pal 5 (Odenkirchen et al., 2001), and integrate its orbit in a static version of one of our TNG potentials at the last snapshot for $13$ Gyr. We remove stars from the cluster given its mass loss rate from equation 2. At each age plotted, we update the upper mass bound based on the current location of the main sequence turn off, which varies slightly with cluster metallicity. For this figure we adopt a metallicity similar to Pal 5 of $\mathrm{[Fe/H]}=-1.56$ (Koch & Côté, 2017), which corresponds to a maximum mass of $0.8\,{\mathrm{M}_{\odot}}$ at present with the MIST isochrones. For the approximate metallicity range relevant to GCs, between 2.5-2.5 and 0.5-0.5 dex, the maximum mass at 13 Gyr ranges from $0.8-0.88\,{\mathrm{M}_{\odot}}$. After several Gyr, the slope of the low-mass end becomes less steep as low-mass stars are preferentially stripped from the cluster. This illustrates that streams realized with individual stars cannot be modeled by simply applying a zero-age IMF, because the cluster evolution significantly changes the mass distribution of the stream, especially for streams which are nearly or completely disrupted.
We assign stream particles numerically according to Equations 6 and 7. At each timestep, we take the mass-loss rate of the cluster and calculate how much mass needs to be removed. Then we remove 2 stars at a time (one for the leading tidal tail and one for the trailing, without replacement) from the bound stellar pool using the same method as described above, until the total mass released at a given timestep equals the appropriate amount according to the cluster mass-loss rate. This is repeated until the present day, giving us $N_{\rm stars}(t)$.
2.3.3 Orbit Integration
Given the number of stars released at each timestep, we integrate the stream orbits from when the stars are released (up to $\sim 3.5$ Gyr ago) until the present day. Using the Chen et al. (2025b) particle-spray method, stars are released near the Lagrange points, with their exact phase-space coordinates sampled from a multivariate gaussian distribution. Each cluster orbit is integrated in the combined potential of the galaxy and the cluster’s self-gravity, which is modeled as a Plummer sphere (Plummer, 1911). We continuously decrease the mass of the Plummer sphere according to the evolving mass of the cluster. Stars are released until the present day or until the cluster is fully disrupted, whichever occurs first. Once every stellar orbit has been completed we are left with the final stream resolved with individual stars at the present day. Repeating this process for each GC gives us the final stream catalog for each halo.
Figure 3: Four example streams with stars color-coded by their ejection lookback time. Each stream’s present day galactocentric radius is shown down and to the left of its progenitor location, as indicated by region where the ejection time approaches 0 Gyr.
Figure 4: Examples of streams from the catalog at a range of present day galactocentric radii with a photometric cut $g<24.5$. The top stream subtends the longest extent on the sky at nearly $250^{\circ}$ and its close proximity to the Galactic center has heated it and caused the stream to extend in $\phi_{2}$. The middle stream progenitor fully disrupted approximately $470\rm\penalty 10000\ Myr$ ago, causing a gap along the stream profile. The final stream at $47$ kpc from the Galactic center is dynamically the coldest, making it the shortest and thinnest of the three shown.
3 Results
In this section we calculate overall statistics for our catalog of disrupted GCs and make photometric predictions for present and upcoming missions. For plots we adopt the astropy convention and assume the Solar System to have galactocentric coordinates of $(-8\penalty 10000\ \mathrm{kpc},0,20.8\penalty 10000\ \mathrm{pc})$ in a right-handed system. Unless otherwise stated, we display only stars with an apparent magnitude $g<24.5$, the limit of a single visit for LSST (LSST Science Collaboration et al., 2009). All our photometry calculations are done with MIST Isochrones (Choi et al., 2016; Dotter, 2016). We also note that for our photometric predictions, we do not include extinction effects, and we perform our analysis across the entire sky.
3.1 Catalog Overview and Statistics
Our model produces a full catalog of the tidal remnants of GCs that had non-zero mass at a lookback time $\sim$3.5 Gyr. The catalog provides data for each stream and their progenitor GCs in two data files. The first file contains masses, positions, velocities, energies, and ejection lookback times in the galactocentric frame for each star at the present time, for each stream. The second file contains information on the progenitor GCs, including their mass history during the period of stream generation, the positions and velocities of their orbits, the tidal tensor along their orbits over the period of stream generation, the present-day age, and $\mathrm{[Fe/H]}$ for each cluster. Our repository on Github [^1] https://github.com/cholm-hansen/StreamCatalogs contains a description of the catalog, provides instructions for downloading it, and contains an example notebook showing how to calculate properties of the streams.
In total, our four halos (523889, 519311, m12i, m12w) contain 287, 401, 391, and 392 streams respectively. Figure 2 shows every stream for one of the TNG halos (523889) plotted on the sky with our fiducial magnitude cut. Each remnant is colored by its metallicity, which is the same value for each star of a given cluster. There is rich variation in the structure of tidal features across the entire population. Toward the Galactic center, there is a spherical grouping of stars from multiple clusters similar in morphology to the Galactic bulge (Wegg & Gerhard, 2013). These are predominantly clusters that have very small pericenters, where the tidal field is too strong to allow long, thin streams. At higher Galactic latitudes, there is an approximately uniform distribution of stars which have deviated significantly from their progenitor, identifiable as members of the stellar halo. Finally, there are many long and thin features which are clearly identifiable as stellar streams.
Figure 5: Variation of each stream in $(\phi_{1},\phi_{2})$ coordinates as a function of the galactocentric radius of the progenitor orbit for stars with $g<24.5$. The top row shows the dispersion for each stream in both $\phi_{1}$ and $\phi_{2}$, where as the bottom row shows the dispersion in the proper motions of each coordinate. Filled circles show thin streams, while open circles show wide streams. Streams closer to the Galactic center are dynamically hotter than streams further away and therefore exhibit a larger variation. Lines show best-fit power law relations at $r>10$ kpc for the three magnitude cuts: LSST (solid line), Roman (dot line), and Gaia (dot-dashed line).
Not every cluster’s tidal feature has a morphology that would be traditionally identified by an observer as a stream. We investigated distributions of the width and axis ratios of all features in the catalog to see if there was an obvious way to group them into ’classical’ streams with more characteristic thin tidal tails and wider streams. For this purpose, for each GC we first rotate its observed tidal debris on the sky into a frame aligned with the two principle components of the escaped stars, denoted as $(\phi_{1},\phi_{2})$ in angular sky coordinates. This is done by first constructing a rotation matrix $R$ via the eigenvalues and eigenvectors of the covariance matrix of star positions $(x,y,z)$. We then rotate the stars into this new frame $(x^{\prime},y^{\prime},z^{\prime})$, where for a thin stream, $x^{\prime}$ is orthogonal to the long axis of the stream, $y^{\prime}$ is aligned with the long axis, and $z^{\prime}$ is orthogonal to the short axis. $\phi_{1}$ is then the longitude angular coordinate (length) and $\phi_{2}$ is the latitude angular coordinate (width) in the stream-aligned frame. We find that inspection of the stream population in $(\phi_{1},\phi_{2})$ coordinates does not reveal evidence of two distinct populations. In light of this, and the fact that there is no consensus definition of a stream in the literature, we use the term ’stream’ for all objects in the catalog but we separate them into thin and wide streams below.
The average combined mass of all streams in each halo is $2\times 10^{7}\,{\mathrm{M}_{\odot}}$, ranging from $1.6\times 10^{7}\,{\mathrm{M}_{\odot}}$ (m12w) to $2.5\times 10^{7}\,{\mathrm{M}_{\odot}}$ (519311). The average total mass of stars in features within $3$ kpc of the Galactic center, roughly corresponding to the radius of the Galactic bulge, is $1.3\times 10^{7}\,{\mathrm{M}_{\odot}}$, ranging from $0.93\times 10^{7}\,{\mathrm{M}_{\odot}}$ (m12w) to $1.8\times 10^{7}\,{\mathrm{M}_{\odot}}$ (519311). Figure 3 shows four example thin streams from halo 523889, which span a range of present-day galactocentric radii 5205-20 kpc. Stars are colored by how long ago they were released from the cluster. A different set of streams spanning 5505-50 kpc is shown in Figure 4. For each stream in this Figure, the progenitor location is placed at $\phi_{1}=0^{\circ}$. The innermost stream, at a present day galactocentric radius of $5.4$ kpc, is the widest and has noticeable width variations along its track. The progenitor GC of the stream in the middle panel is fully disrupted at the present day, causing a gap to have formed in its center. There is also a slight asymmetry in the tidal tails caused by the projection into the $\phi_{1}-\phi_{2}$ frame. The outermost stream, at $47$ kpc, is the shortest, thinnest, and least massive, being subjected to the weakest tidal field out of the three. Collectively, these Figures illustrate the wide range of features and morphologies present in the overall MW stream population.
Figure 6: Cumulative distribution of the stream width-to-length ratio with two upper limits on the width: $\sigma_{\phi_{2}}<5^{\circ}$ and $\sigma_{\phi_{2}}<10^{\circ}$. We adopt $\sigma_{\phi_{2}}/\sigma_{\phi_{1}}=0.4$ based on the flattening of the CDF at that location to differentiate thin streams from wide streams.
3.2 Thin vs. Wide Streams
Figure 5 shows how standard deviation in $\phi_{1}$, $\phi_{2}$, $\mu_{\phi_{1}}$ [^2] Here and in the rest of the paper, we use $\mu_{\phi_{1}}$ as shorthand for $\mu_{\phi_{1}}\cos{\phi_{2}}$., and $\mu_{\phi_{2}}$ vary as a function of distance to the Galactic center, as would be observed from Earth. Here, $\sigma_{\phi_{1}}$ characterizes the length of the stream, and $\sigma_{\phi_{2}}$ characterizes its width. These statistics are calculated for the stars in each stream that are below our fiducial magnitude cut, $g<24.5$.
We distinguish here between what we consider wide streams (open circles) and thin streams (filled circles). These two groups are classified in the following way: first, we rotate each stream into galactocentric $(\phi_{1},\phi_{2})$ coordinates. This is done to emphasize the heating caused by an orbit being close to the Galactic center, and to de-emphasize observation effects, such as a stream with a blob-like morphology which appears thin due to its distance from the Earth. Any stream with $\sigma_{\phi_{2}}$ exceeding $10^{\circ}$ is considered a wide stream, which captures features that are too thick to be traditionally classified as a thin stream. Then we make a cumulative distribution of the ratio of a stream’s width to its length, $\sigma_{\phi_{2}}/\sigma_{\phi_{1}}$, in Figure 6. Based on the flattening of the CDF gradient, we classify thin streams as having $\sigma_{\phi_{2}}/\sigma_{\phi_{1}}<0.4$, corresponding to axis ratio greater than 2.5:1. We also repeated this procedure with an initial cut of $\sigma_{\phi_{s}}$ exceeding $5^{\circ}$ to check if our choice of initial cut significantly affects the CDF. As can be seen from Figure 6, there is a minimal difference between an initial cut of 5 or 10 degrees; for our classification we use 10 degrees.
Beyond the inner galaxy, these relations all appear to follow a power law, but with significant intrinsic scatter. For each quantity $X$, we use maximum likelihood estimation to calculate the best-fitting power law relation of the form $X\propto r^{\alpha}$ for all points beyond $r>10$ kpc, because at smaller radii the relation visually appears to break down. This is influenced by the solar radius being at $\approx 8$ kpc, but we also find that the relation breaks around the same radius when the quantities in Figure 5 are calculated as seen from the Galactic center. To estimate the uncertainties for each slope, we resample our data $1000$ times and use the bootstrapping technique. For the standard deviation in $\phi_{1}$ and $\phi_{2}$ (in the top row of Figure 5) the best-fit slopes are $\alpha=-1.8\pm 0.3$ and $\alpha=-2.0\pm 0.5$. For the dispersion in proper motions along $\phi_{1}$ and $\phi_{2}$ (the bottom row of the figure), the slopes are $\alpha=-3.1\pm 0.6$ and $\alpha=-3.2\pm 0.5$.
We also recompute what the slopes would be using magnitude cuts for Roman (dashed line) and Gaia (dot-dashed line). Details of how the cuts are made are in the next section (3.3). Overall the fits are mostly consistent between the three instruments. The width fit with the Gaia magnitude cut shows the greatest deviation from the other two, producing a slightly steeper ($\alpha=-2.2$) best-fit slope. This can be explained from the fact that the Gaia samples are most incomplete at large galactocentric radii. This suggests that width estimates for GC streams with members stars confirmed with Gaia may be underestimated.
Figure 7: Cumulative distribution of the number of observable stars in each stream according to photometric completeness limits for Gaia ($G<19$), LSST ($g<24.5$), and Roman ($z<27.5$). Based on their reported magnitude limits, LSST and Roman will both drastically expand our ability to identify MW stellar streams.
3.3 Stream Photometric Predictions
[…]
We integrate our PDFs to get cumulative distributions of the number of streams that contain $N_{\rm stars}$ observable stars for a given instrument, for a given halo: $N(>N_{\rm stars})$. This result is shown in Figure 7. For each instrument, the region plotted encapsulates all four individual distributions for each halo. The uncertainty arises from the differing number of total streams in each halo. We also plot the theoretical limiting distribution if every stream star in the catalog were observable. The figure shows how impactful LSST and Roman will be in identifying new streams. For example, Gaia is only able to observe at least $10^{4}$ stars in about $60$ streams, but LSST and Roman would be able to observe the same number of stars in hundreds of streams. Roman would be able to detect at least 1000 stars in almost the entire sample across all four halos.
In Figure 8 we divide the streams into three approximately equal sized-groups based on their present-day galactocentric radius, for the adopted LSST magnitude limit of $g<24.5$. Even for streams beyond 12 kpc, LSST will have resolving power to identify several dozen, up to potentially more than one hundred member stars according to our catalog, depending on the true intrinsic radial distribution of streams in the outer halo. An extended catalog of identified GC stellar streams at large galactocentric radii will be crucial for better understanding the difference between baryonic and DM perturbations to streams, which will be key if streams are to one day probe the subhalo mass function and constrain the nature of DM.
Figure 8: Number of observable stars per stream binned by their present-day galactocentric radii. Each shaded region encapsulates the range between all four halos. For example, the number of streams beyond 12 kpc with 100 or more observable stars ranges from 6018060-180 depending on the halo. The uncertainty arises primarily from the varying number of streams in each halo. Depending on the true number, our catalogs predict there are anywhere from $\sim$60 to well over one hundred streams beyond 12 kpc with 1000 or more stars below the detection limit of LSST.
Table 1: The 25%-75% percentile range of the lowest observable stellar mass (in $M_{\odot}$) for streams in a given bin of galactocentric radius, averaged across streams in all four halos.
Radial range $0-4\rm\penalty 10000\ kpc$ $4-12\penalty 10000\ \rm kpc$ $12\penalty 10000\ \rm kpc$
Gaia $(G<19)$ 0.640.770.64-0.77 0.350.700.35-0.70 0.750.820.75-0.82
LSST $(g<24.5)$ 0.200.360.20-0.36 0.110.250.11-0.25 0.250.610.25-0.61
Roman $(z<27.5)$ $0.1$ $0.1$ 0.10.200.1-0.20
Table 1 lists the lowest mass of stream stars detectable by the three instruments in each bin of galactocentric radius from Figure 8. For each stream in each halo, we compute the lowest mass observable according to each instrument’s limiting magnitude, then for each bin we compute the 25th and 75th percentile values. The table displays these values averaged across the four halos. Stars in the middle bin are typically at the closest heliocentric distances, which will allow LSST and Roman to sample the mass function even below $0.2\,{\mathrm{M}_{\odot}}$. Gaia is also relatively effective in this radius bin, being able to probe down to $\sim 0.3\,{\mathrm{M}_{\odot}}$ for a quarter of the streams within this galactocentric distance range. However, in the outer bin, the Gaia limit increases dramatically. The percentiles reported here are only for the streams in which there is more than one star below the detection limit of Gaia; in every halo, there are streams that Gaia misses entirely. The exact number varies according to the varying quantity and radial distribution of streams in each halo, from a minimum of 4 in 523889, to a maximum of 91 in 519311. The FIRE halo m12w, which most closely matches the MW in terms of its GC radial distribution (see Appendix C), there are 19 streams for which Gaia can observe no stars. Given the fact that at minimum several dozen stars are needed to identify a stream with current methods (e.g., Malhan & Ibata, 2018), this result suggests that there are anywhere from several dozen up to over one hundred streams in the MW halo undetectable with Gaia.
The outer streams, at a median radius of 28.5 kpc, should still reveal stars all the way down to $0.2\,{\mathrm{M}_{\odot}}$ with Roman. This means that Roman has the potential to provide an essentially complete view of all disrupted GCs in the Milky Way.
4 Discussion
Our model catalog of 1473 tidal features from disrupted GCs across four cosmological halos contains a diverse set of objects including hundreds of thin streams, which can be used for many purposes such as analysis of a stream population in a MW-like galaxy, mock photometry of individual streams, modeling of streams to constrain the Galactic potential, and testing stream detection algorithms. Here we describe additional possible uses of the catalog, analyze our initial results further, and discuss possible limitations.
4.1 Orbital Properties of Thin and Wide Streams
Given the orbital information of the GC progenitors for the last several Gyr, we can further distinguish between the types of orbits that form long, thin stellar streams and those that do not. In Figure 9 we show the energy-angular momentum diagram for all 287 progenitor orbits for the features in halo 523889. In our coordinate system prograde orbits are on the right (positive $L_{z}$) and retrograde orbits are on the left (negative $L_{z}$). The majority of thin streams based on our definition have energies greater than $0.1\penalty 10000\ \rm kpc^{2}\penalty 10000\ Myr^{-2}$, indicating that orbits more strongly bound than this (thereby residing closer to the Galactic center) are heated to the point where they do not form thin coherent streams. However, there are also many streams on prograde circular orbits which remain thin deeper into the potential well. Notably, the current known phase-space distribution of GC streams in the MW (Ibata et al., 2024) does not have such streams. It is possible that these streams have evaded detection thus far; they are closer to the Galactic center and appear on the sky in regions of higher background stellar density, which makes them harder to detect. Further analysis of the orbits in connection with the resulting tidal features should help to clarify what types of streams to expect on what types of orbits.
Figure 9: Characterization of orbits for all objects in 523889 in the space of the specific energy $E$ and the $z$ component of angular momentum. Tidal debris which resides deeper in the Galactic potential tend to not form thin streams. The solid line indicates the limit of a prograde circular orbit, while the dashed line represents the limit of a retrograde circular orbit.
4.2 Impact of Future Observatories
Our findings agree with previous results that both LSST and Roman will greatly expand our knowledge of streams in the MW and other galaxies in the Local Group (Pearson et al., 2019; Shipp et al., 2023; Aganze et al., 2024; Pearson et al., 2024). Specifically, we find that the number of observable stars per stream will increase drastically.
We can also use these data to estimate the number of streams potentially observable at present that have evaded detection. Using our adopted definition of a thin stream, there are between 60 and 124 thin streams with 100 or more observable stars with Gaia in each halo, compared with the 80 or so known GC stellar streams. This suggests the possibility of several dozen missing streams within the detection limit of Gaia. However, the number may even be greater than this, since we are only tracking streams which formed in the last $3.5$ Gyr (see Pearson et al. 2024). One possible reason for evading detection is extinction from MW dust near the Galactic plane. Out of our set of thin streams with at least 100 stars below the Gaia detection limit, the total number with progenitor locations on the sky with $|b|<10^{\circ}$ ranges between 36 – 64 for our four different halos. These streams would be partially or completely obscured in optical wavelengths due to foreground dust. A future infrared survey astrometry telescope could make it possible to detect such streams, such as the proposed GaiaNIR mission (Hobbs & Høg, 2018).
4.3 Synthetic Photometry of Individual Streams
It is also possible to analyze individual streams in the catalog and create analogs to real observed systems. An example is in Figure 10, which shows the color-magnitude diagram (CMD) of a stream with similar mass and orbit to Pal 5. Its progenitor has an orbital pericenter of $11$ kpc, apocenter of $21$ kpc, and a present-day mass of $1.6\times 10^{4}\,{\mathrm{M}_{\odot}}$. In order to match the observed properties of Pal 5, for this Figure only we keep the Sun fixed at $8.122$ kpc from the Galactic center, but perform a rotation in the xyx-y plane to best match Pal 5’s radial velocity ($58.6\penalty 10000\ \rm km\penalty 10000\ s^{-1}$; Baumgardt et al., 2019) and heliocentric distance ($21.9$ kpc; Baumgardt & Vasiliev, 2021). The optimal rotation gives our mock cluster a heliocentric distance of $20.9$ kpc and $r_{v}=-64.8\penalty 10000\ \rm km\penalty 10000\ s^{-1}$. We use the MIST isochrone for an age of $10.355$ Gyr and metallicity of $\mathrm{[Fe/H]}=-1.56$ to compute mock SDSS $g$ and $r$ magnitudes for the stars with their updated distances, assuming no extinction along the line of sight.
Stars along the stream have slightly different heliocentric distances than the progenitor, which creates measurable scatter around the original isochrone. We treat the mock CMD as if it were observed data and compute the best-fitting isochrone for all stars with $g<24.5$ to the widened CMD. We sample a $100\times 100$ grid of ages and metallicities ranging from $\{9,13\}\penalty 10000\ \rm Gyr$ and $\{-2,-0.5\}\penalty 10000\ \rm dex$, and minimize the summed squared distance from each point to the isochrone. The best fitting isochrone from this search has an age of $9.2\penalty 10000\ \rm Gyr$ and $\rm Fe/H]=-1.23$. The two curves are similar near the main sequence turnoff but begin to deviate by about 0.1 dex at $r\approx 25$, where the stars were not included in the fitting. There is also a notable discrepancy on the red giant branch where there are fewer stars. Notably, if all stars are included, we get a much closer fit giving an age of $9.97$ Gyr and $\rm[Fe/H]=-1.49$. This example shows that deeper imaging, such as that of Roman, will allow for significantly better constraints on a stream progenitor’s age and metallicity given its CMD.
We can also use this mock CMD to provide a practical guide for future observations trying to select a relatively complete sample of stream stars for follow-up spectroscopy, while minimizing contamination from the stellar background. We find that $80\%$ of the stream stars are located within $\pm 0.1$ dex in color from the best-fit isochrone for stars included in the original fit. Filtering stars based on the spread from an isochrone is commonly used in identifying stream candidates (e.g., Valluri et al., 2025), and our analysis shows that it should indeed capture the majority of the stream stars.
len
Figure 10: Color-magnitude diagram for a stream with similar apocenter and pericenter values to Pal 5, color-coded by stellar mass. The corresponding cluster in the catalog has an age of $10.36$ Gyr and we set the metallicity to $\mathrm{[Fe/H]}=-1.56$. The MIST isochrone for a stellar population with this age and metallicity at a distance of $20.9$ kpc is shown as a dashed line. Next to it is a solid line, representing the best-fit isochrone from the data with $g<24.5$ ($9.2$ Gyr, $\mathrm{[Fe/H]}=-1.23$). The scatter of stars is due to stream stars having different distances from Earth. The dotted lines located at $\pm 0.1$ dex in color from the best-fit isochrone show the bounds needed to capture 80% of all stream members.
4.4 Connection to Host Potential
While not done explicitly here, the catalog along with its provided potential can be used to better understand the connection between a stellar stream population and its host galaxy. Primarily, this involves using streams to measure the enclosed mass profile of the galaxy, which can then be applied to real streams in the MW. This can also be extended to measuring the shape of the MW halo, which has been done in previous work but has produced contrasting results depending on the assumed model parameters. More work is needed along these directions, and our catalog serves as a possible useful tool.
Another promise of streams is studying the connection between the stream population and the assembly history of the host halo (Carlberg, 2017; Bonaca et al., 2021; Li et al., 2022). Our GC formation and stream models are based on cosmologically-formed halos, for which we know the true assembly history. Such a system has proved valuable for reconstruction of the galactic assembly histories with GCs (Chen & Gnedin, 2024b). With our stream catalog, a more comprehensive analysis can be applied to better understand the MW assembly history.
4.5 Scaling of Stream Length
In Section 3 we found that the length of the stream population follows a power law dependence on galactocentric radius, $\sigma_{\phi_{1}}\propto r^{-1.8\pm 0.3}$. The idea that the length distribution should follow a power law is not immediately obvious. To understand this result better, we consider an idealized model of a progenitor GC with mass $m$ on a circular orbit at radius $r$ with orbital frequency $\Omega$. We assume that the enclosed mass of the Galaxy follows the relation $M(r)\propto r^{1-\gamma}$ with $\gamma\approx 0$ at the range of radii 1010010-100 kpc, where we measure the power-law relation. For the real Milky Way, Shen et al. (2022) estimate $\gamma\approx 0.4$; for our simulated galaxies, we estimate an average $\gamma\approx 0.2$. Johnston (1998), Johnston et al. (2001), and Amorisco (2015) showed that the length of the stream is expected to grow with time as
$\sigma_{\phi_{1}}\sim t\,r_{t}\,\frac{\partial\kappa}{\partial r}$ (8)
where $\kappa$ is the epicycle frequency and $r_{t}$ is the tidal radius. For our adopted mass distribution $\kappa=(2-\gamma)^{1/2}\,\Omega$. The instantaneous Lagrange points define the tidal radius as
$r_{t}\approx\left(\frac{m}{2M(r)}\right)^{\frac{1}{3}}r$ (9)
however for eccentric GC orbits the tidal radius may be set closer to the peri-center $r_{p}$; we consider this possibility below. For the circular orbit we obtain
$\sigma_{\phi_{1}}\propto r^{-\frac{4}{3}-\frac{\gamma}{6}}\,t$ (10)
and therefore, for a population of streams whose duration times are comparable, in this simple model we would expect the lengths of streams to follow a power law, albeit with a shallower slope ($\approx-1.4$) than the one found for our catalog.
One possible reason that the measured slope (1.8-1.8) would be steeper is that GCs at large galactocentric radii tend to be on highly radial orbits, for which effective tidal truncation happens at smaller tidal radii than given by Equation 9. Additionally, streams on radial orbits spend a longer duration near apocenter, where the tidal tails tend to compress as the stream stars cluster more tightly together in action space.
If we include only streams with eccentricities less than $0.4$ (of which there are 59 beyond 10 kpc) to select orbits closer to circular, we find that the best-fit slope becomes $\alpha=-1.7\pm 0.3$, which is indeed shallower than the result with all streams but still in tension with our toy model. A more robust calculation of the length scaling would take into account non-circular orbits, as well as the expansion and compression of streams near pericenter and apocenter. We do not attempt this here, but nevertheless we find this simple model useful in trying to understand why the lengths should follow a power law relation.
4.6 Modeling Limitations
Selecting galaxies from cosmological simulations for our model calculation leads to two numerical limitations. First, we are limited by both the spatial and temporal resolution of the simulations. Since our model GC initial positions are assigned to particles in TNG and FIRE, they can be limited by each simulation’s force-softening length. In FIRE, this is only $4$ pc for stellar particles, but for TNG it is $\sim 300$ pc, which can lead to errors in the distribution of GCs in the inner galaxy at a lookback time of $3.5$ Gyr. Moreover, the time resolution between snapshots used in the potential fitting varies between 100 and 200 Myr for TNG and 10 Myr - 1 Gyr for FIRE, which washes out important baryonic components of the potential that can impact streams, such as the spiral arms and the Galactic bar. There is also no LMC analog in our selected halos, which has been shown to leave significant perturbations on streams in its vicinity (e.g., Brooks et al., 2025). Higher resolution large-scale simulations will be needed to model all of these affects more accurately and study their impact on the total Galactic stream population.
Second, the fitting of the potential using BFE has its own limitations when considering the appropriate choices of $l_{\rm max}$ and $m_{\rm max}$. Selecting too low a value for these parameters, particularly $l_{\rm max}$, will smooth out too much of the halo’s features and substructure. However, too high a value of $l_{\rm max}$ will cause the potential to be overfit in regions where the particle numbers are low. This is especially true close to the Galactic center, where the number of particles enclosed within a spherical shell becomes small. For example, out of all the particles present at the last snapshot for TNG halo 523889, only $3\%$ are within 1 kpc of the Galactic center, and only $11\%$ are within 4 kpc, which roughly corresponds to the scale length of the MW disk. At the same time, it is important to model the potential accurately at small galactocentric radii, since the cluster mass-loss on eccentric orbits is dominated at pericenter. We made the selection of $l_{\rm max}$ and $m_{\rm max}$ after careful consideration of both competing issues (see Appendix A), but further improvements in this area will be essential for high-precision modeling of streams in time-evolving potentials.
5 Conclusions
Here, we summarize the main takeaways of this work:
  • We have generated a catalog of GC stellar streams in four MW-like galaxies within a cosmological context of the TNG and FIRE simulations. The catalog combines the Chen & Gnedin (2024a) GC evolution model with the Chen et al. (2025b) algorithm to generate streams in the time-evolving potentials. The catalog contains the full phase-space information for individual stars in each stream, as well as orbital history and physical properties for stream progenitors.
  • We find that the model stream length, width, and proper motions in the MW halo follow power law scalings with galactocentric radius. The best-fit slopes for the length and width are $\alpha=-1.8$ and $\alpha=-2.0$, respectively.
  • We apply mock photometry to our streams and quantify the number of observable stars per stream with Gaia, LSST, and Roman. We find that both LSST and Roman should discover hundreds of new streams in the Milky Way. In particular, Roman will be able to observe stars down to $0.1\,{\mathrm{M}_{\odot}}$ for the majority of streams.
  • We make our catalog publicly available as a tool for studying GC streams in MW-like galaxies.
The repository for the catalogs is hosted at https://github.com/cholm-hansen/StreamCatalogs, which contains a description of the catalog and its components, instructions for downloading, and an example notebook showing how to use and perform calculations with it.
Acknowledgments
The authors thank Monica Valluri, Eric Bell, Eugene Vasiliev, Neil Ash, and Brigette Vazquez Segovia for insightful discussions. CHH, YC, and OYG were supported in part by National Aeronautics and Space Administration through contract NAS5-26555 for Space Telescope Science Institute programs HST-AR-16614 and JWST-GO-03433. This research benefited from the Gravity in the Local Group conference hosted by the McWilliams Center for Cosmology and Astrophysics, Carnegie Mellon University. This research was supported in part through computational resources and services provided by Advanced Research Computing at the University of Michigan, Ann Arbor.
Software
agama (Vasiliev, 2019), astropy (Astropy Collaboration et al., 2022), isochrones (Choi et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020)
Appendix A Calibration of the Galactic Potential
Figure 11: Convergence of the halo and disk potentials for TNG galaxy 523889. For each data point, we increment $l_{\rm max}$ or $m_{\rm max}$ and plot the average absolute relative difference across the initial cluster locations compared to the previous point. The solid lines show the convergence using the first TNG snapshot used in our model and the dotted lines shows the last snapshot.
Figure 12: An example of the effective tidal frequency (Eq. 3) along the orbit of a model cluster with similar apocenter and pericenter to Pal 5. The line shows the analytic calculation of the tidal tensor using the BFE potential in agama. Each spike corresponds to a tidal shock near the orbital pericenter. The vertical dashed lines indicate the epochs when full TNG snapshots are available to calculate the tidal tensor directly from the baseline simulation. The only four available data points emphasize the advantage of using the agama orbits to compute $\Omega_{\mathrm{tid}}$.
From a lookback time of $\sim$3.5 Gyr until the present, we use potentials fit with BFEs in agama to calculate orbits of progenitor GCs and tidally stripped stars, as well as for the calculation of the mass-loss rate of the GCs along their orbit. A two-component BFE is fit to simulation snapshots, which is then linearly interpolated in time to create the time-dependent potential. The DM particles are fit with a multipole expansion:
$\Phi(r,\theta,\phi)=\sum_{l=0}^{l_{\mathrm{max}}}\sum_{m=-l}^{l}\Phi_{l,m}(r)Y_{l}^{m}(\theta,\phi)$ (11)
where $Y_{l}^{m}(\theta,\phi)$ are the usual spherical harmonics. The star and gas particles are modeled with an azimuthal Fourier harmonic expansion in $\phi$:
$\Phi(R,z,\phi)=\sum_{-m_{\mathrm{max}}}^{m_{\mathrm{max}}}\Phi_{m}(R,z)e^{im\phi}$ (12)
Figure 13: CDFs of the mass, metallicity, and radial distribution of surviving GCs in our catalogs at the present day. The MW is plotted for comparison. Our GCs are in good agreement with the present day mass and radial distribution of MW GCs, which are the most important parameters for generation of stellar streams.
Before the potentials are fit, we rotate the particles in each snapshot to a common reference frame aligned with the principal axis of the disk particles at the last snapshot. This is needed for proper fitting of the Fourier expansion terms, in addition to needing one consistent coordinate system in order to interpolate between adjacent snapshots. While the disk orientation does change slightly between the first and last snapshot, we find that for three of the four halos used, this change is only a few degrees, which does not have a significant impact on the fit. For the one with a significant change in the disk orientation (m12i), we opt for a BFE using only multipole expansion, as described in Section 2.2.
For both the halo and disk potentials, we sample coefficients $\Phi_{l,m}(r)$ and $\Phi_{m}(R,z)$ on logarithmically spaced grids and spline interpolate between the individual points to get the final analytic potential at each snapshot. The time-dependent potential is then constructed by linearly interpolating between the individual snapshot potentials. For the multipole expansion, no symmetry is assumed and all $m$ terms are kept for each corresponding $l$. To determine the highest order of $l$ in the multipole expansion and $m$ in the Fourier harmonic expansion, we perform a convergence test by evaluating the change in each respective potential at the locations of the GCs 3.5 Gyr ago for one of our halos (523889). For the disk potential, we start at $m_{\rm max}=1$ and compute the relative absolute change in potential between $m_{\rm max}-1$ and $m_{\rm max}$, averaged at the locations of all the GCs. The same is done for the halo potential, except we increment $l_{\max}$ by 2 instead of 1. The result is shown in Figure 11. The shorter lines are for $m_{\rm max}$ and the longer ones are for $l_{\rm max}$. Additionally, the solid lines are for the first snapshot used for the BFEs ($z=0.3$), and the dashed lines are for the last snapshot ($z=0$). Based on the convergence of both the disk and halo potentials, we adopt $l_{\rm max}=12$ and $m_{\rm max}=6$.
Appendix B Comparison of Low and High Time Resolution for Mass-Loss Calculation
The fitting of the BFE potentials allows us to calculate the mass-loss rate along the GC orbits at higher resolution than is available with the GC evolution model in Chen & Gnedin (2024a). In Figure 12, we show our computed effective tidal frequency (defined from the eigenvalues of the tidal tensor), for a selected cluster in the catalog on a similar orbit to Pal 5. Each peak in $\Omega_{\rm tid}$ is from a pericenter passage of the GC on its orbit. By comparison, the Chen & Gnedin (2024a) model only computes $\Omega_{\rm tid}$ at full TNG snapshots, the timestamps of which are shown with vertical dashed lines. For our higher resolution calculation using BFE, agama interpolates the computed solution to give us the values of $\Omega_{\rm tid}$ on $10^{3}$ uniformly space intervals within the last $\sim$3.5 Gyr. Our higher resolution calculation is used to better capture the mass-loss of the GCs along their orbits.
Appendix C Properties of Surviving GC Population
For the creation of our GC stream catalogs, we modified the GC evolution model in Chen & Gnedin (2024a) to output cluster masses at each snapshot and subsequently had to re-run the model on the four cosmological halos. Due to inherent random number generation in the code, our GC populations are slightly different than the published catalogs in Chen & Gnedin (2024a). Here we show the properties of each GC population in comparison with the MW. Figure 13 shows the cumulative probability distribution functions (CDFs) for the mass, metallicity, and radial distributions of surviving GCs in our catalog, with masses greater than $10^{4}\,{\mathrm{M}_{\odot}}$ in the present day. Observed CDFs for the MW GC population are displayed with a thicker line. Our GC populations are matched well particularly to the mass and radial distributions of the MW GCs, both of which are more important for the purposes of generating stellar streams.